OpenStudy (anonymous):

(5a^2/3)(4a^3/2) simplify and evaluate where appropriate (exponents or logarithms)

OpenStudy (anonymous):

Is it a^(2 / 3) and a^(3 / 2) or (a^2) / 3 and (a^3) / 2

OpenStudy (anonymous):

the first one

OpenStudy (anonymous):

If it's the first case, use the law a^m • a^n = a^(m + n)

OpenStudy (anonymous):

I don't know what that law is...

OpenStudy (jdoe0001):

keep in mind that \(\huge a^{\frac{n}{m}} = \sqrt[m]{a^n}\)

OpenStudy (anonymous):

ohhhhhh. thank you!!